### v2

Abstract
This paper explore the differential geometric structure of statistical manifolds defined by Gaussian distributions, focusing on their representation in a dual coordinate system. Using Amari’s information geometry framework, we derive explicit expressions for the Fisher information metric and compute the associated torsion and curvature tensors in the dual setting. Our results show that, unlike the standard coordinate system where curvature and torsion vanish, the dual system exhibits non zero values for these three geometric quantities. This highlights how dual parameterizations can affect the intrinsic geometry of statistical models.
keywords:
Tensor curvature, Riemannian metric, Torsion.
1 Introduction
Information geometry is a powerful mathematical framework that applies the methods of differential geometry to the study of statistical models and probability distributions. By treating families of probability distributions as points on a manifold, this field provides a geometric interpretation of key statistical concepts such as divergence, estimation, and entropy. Among its central tools are the Fisher information metric and dual affine connections, which allow one to study geometric invariants like curvature and torsion within a statistical setting.
The foundational work of C.R. Rao [2], first demonstrated the relevance of differential geometry in statistics by introducing a Riemannian metric defined through the Fisher information matrix, enabling the computation of geodesic distances between distributions. Building on this insight, S. Amari and H. Nagaoka [2], developed a comprehensive theory of information geometry, formalizing the notions of dual coordinate systems and affine connections on statistical manifolds. Their theory shows that many statistical models, including the Gaussian family, exhibit flatness and torsion free properties in their natural (exponential or mixture) coordinate systems. However, the behavior of geometric properties under dual coordinate transformations remains a subtle and open question. While it is well known that the Gaussian manifold is flat and torsion-free in its natural coordinate system $\theta$, it is not always clear whether such nonlinear geometric invariants are preserved when one moves to a dual coordinate system, such as $\xi=(\mu,\mu^{2}+\sigma^{2})$? This leads to the central problematic of this work: To what extent does the dual coordinate transformation affect the intrinsic nonlinear geometry (curvature and torsion) of a Gaussian statistical manifold? Are these geometric invariants preserved, distorted, or enriched in the dual setting? Several researchers have addressed related questions. M.N. Boyom [1, 3] has explored the geometric and numerical properties of Koszul connections in statistical settings, while J.-L. Koszul, and collaborators [4], have studied deformations of locally flat connections. In more recent works, in [5], investigated the geometric structure of beta distributions within the framework of information geometry. Ovidiu Calin [6], further developed the differential geometric modeling of the Gaussian distribution, providing explicit expressions for curvature and connection coefficients in natural coordinates. In addition, J.M. Lee [7], offered a comprehensive exposition of Riemannian manifolds and curvature tensors that underpins the theoretical foundation of our investigation. It is also worth noting that recent developments have shown the space of vector fields $T^{l}(\theta)$ on a Gaussian manifold can be identified with a submodule of second-degree polynomials with smooth coefficients scaled by $\sigma^{-3}$, i.e., $\sigma^{-3}C^{\infty}(S)[x]^{2}$, offering new algebraic insights into the tangent bundle structure. In this work, we extend these investigations by focusing on the Riemannian and affine geometric structures of the Gaussian statistical manifold when expressed in the dual coordinate system $\xi=(\mu,\mu^{2}+\sigma^{2})$. We derive explicit formulas for the Fisher information metric, the Christoffel symbols, the torsion tensor, and the Riemann curvature tensor in this dual system. We prove that the curvature in the $\xi$-coordinate system is equal to
$\mathcal{K}$ $=$ $\frac{16\mu^{2}\sigma^{8}+12\mu^{4}+6\sigma^{10}-\mu\sigma^{8}-2\mu^{2}\sigma^{2}+10\mu^{3}\sigma^{6}+92\mu^{3}\sigma^{8}+48\mu^{5}\sigma^{6}}{4\sigma^{10}}$
$+\frac{12\mu\sigma^{10}+24\mu^{5}-6\mu^{3}}{4\sigma^{10}}$
A key outcome of our analysis is the emergence of a non zero scalar curvature, which contrasts with the flatness observed in the natural coordinates. This result suggests that dual parametrization may introduce richer nonlinear geometric features, with implications for the stability and global structure of statistical models. The structure of the paper is as follows: Section 2 reviews the basic theoretical framework of statistical manifolds and introduces the Gaussian model. Section 3 presents the derivation of the Riemannian metric in the dual $\xi$-coordinate system. Section 4 is devoted to computing the torsion and curvature tensors in this setting. We conclude with a discussion of the implications of our findings and potential directions for future work, including generalizations to other distributions and geometric invariants.
2 Preliminaries
[…]
2.1 $\theta$-coordinates System and $\xi$-coordinates system..
Let $S$, the statistical manifold
Definition 1.
[2] Let $U$ be an open subset of $S$ and $\phi$ be a homeomorphism of $U$ in $\mathbb{R}^{n}$. Then a point $p\in U$ is a function de $\theta=\left(\theta_{1},\theta_{2},\dots,\theta_{n}\right)\in\mathbb{R}^{n}$ by $\phi$, i.e. $\phi(p)=\theta$ this application $\phi$ is called a function coordinates in the neighbourhood of $p$. It is introduced a coordinate system in $U$ called the Any point $p$ in $U$ is given by the $\theta$-coordinates given by $\theta=(\theta_{i});i\in\{1,2\dots,n\}$.
According to Amari’s[2], the $l$-representation basis is given by $\mathcal{B}_{\theta}l(x,\theta)=\left\{\partial_{1}l(x,\theta),\partial_{2}l(x,\theta)\right\}$ where $\mathcal{B}_{\theta}=\partial_{1},\partial_{2}\}$ is the natural basis. In the case of gaussian model we have $\partial_{1}=\frac{\partial}{\partial\mu}$, and $\partial_{2}=\frac{\partial}{\partial\sigma}$ therefore   $\left\{\frac{\partial}{\partial\mu},\frac{\partial}{\partial\sigma}\right\}=\mathcal{B}_{\theta}$ The base associated to the $l$- representation is therefore $\left\{\partial_{1}l(x,\theta),\partial_{2}l(x,\theta)\right\}$. As
$l(x,\theta)$ $=$ $-\log\sqrt{2\pi}\sigma-\frac{(x-\mu)^{2}}{2\sigma^{2}}$ (1)
It follows that
$\partial_{1}l(x,\theta)$ $=$ $-\frac{1}{\sigma}+\frac{(x-\mu)^{2}}{\sigma^{3}}$
So
$\mathcal{B}_{\theta}l(x,\theta)$ $=$ $\left\{\frac{(x-\mu)}{\sigma^{2}},-\frac{1}{\sigma}+\frac{(x-\mu)^{2}}{\sigma^{3}}\right\}.$
Remark 1.
[2] There are several other coordinate functions which at a point $p$, are such that: $\psi(p)=\xi=\left(\xi_{1},\xi_{2},\dots,\xi_{n}\right)$ where $\psi$ is a homeomorphism of an open $V$ of $S$ in $\mathbb{R}^{n}$ is the $\xi$-coordinate system. The coordinates $\left(\xi_{i}\right),\;i\in\{1,\dots,n\}$, characterize the point $p$ in $\mathbb{R}^{n}$. Let’s give ourselves two system coordinates $\theta$ and $\xi$ There is one and only one correspondence between the coordinates $\theta$ and $\xi$; given the properties enjoyed by the applications $\phi$ and   $\psi$, and the fact that they are homeomorphisms we have the following equations
$\xi$ $=$ $\psi\circ\phi^{-1}(\theta),\textrm{and}\;\theta=\phi\circ\psi^{-1}(\xi)$
The transformation of $\theta$ a would be a diffeomorphism, if the $n$-functions $\xi_{i}\left(\theta_{1},\dots,\theta_{n}\right)$ are differentiable and $Det\left|\frac{\partial\xi_{i}}{\partial\theta_{j}}\right|$ is non-zero.
2.2 Determination of the characteristic elements of the $\xi$- coordinates.
[…]
$T^{l}(\theta)$ $=$ $=\left\{A(x);A(x)=A^{1}\frac{\partial_{1}l(x,\theta)}{\partial\mu}+A^{2}\frac{\partial_{2}l(x,\theta)}{\partial\sigma}\right\}$
and using (2.1) we have
$T^{l}(\theta)$ $=$ $\left\{A(x)\textrm{ tel que }A(x)=A^{1}_{(\mu;\sigma)}(x)\frac{x-\mu}{\sigma^{2}}+A^{2}_{(\mu;\sigma)}(x)\frac{(x-\mu)^{2}}{\sigma^{3}}-\frac{A^{2}_{(\mu;\sigma)}(x)}{\sigma}\right\};A^{1},A^{2}\in C^{\infty}(S)$
So $A(x)\in T^{l}(\theta)$ if and only if there exists $A^{1},A^{2}\in C^{infty}(S)$ such that
$A(x)$ $=$ $\frac{A^{2}}{\sigma^{3}}x^{2}+\left(\frac{A^{1}}{\sigma^{2}}-2\frac{A^{2}\mu}{\sigma^{3}}\right)x+\frac{A^{2}}{\sigma^{3}}\mu^{2}-\frac{A^{1}\mu}{\sigma^{2}}-\frac{A^{2}}{\sigma}$
Let $a=\frac{A^{2}}{\sigma^{3}}$ ; $b=\frac{A^{1}}{\sigma^{2}}-2\frac{A^{2}\mu}{\sigma^{3}}$ et $c=\frac{A^{2}}{\sigma^{3}}\mu^{2}-\frac{A^{1}\mu}{\sigma^{2}}-\frac{A^{2}}{\sigma}$ So $A(x)$ is of the form $ax^{2}+bx+c$ we can be reassured that $a=\frac{A^{2}}{\sigma^{3}}\neq 0$. We have thus prove the following result. Put $frac{1}{\sigma^{3}}C^{\infty}(S)_{2}[x]$ the modulus of polynomials of second degree with coefficients in $\frac{1}{\sigma^{3}}C^{\infty}(S)$. Then $T^{l}(\theta)$ is a subset of $\frac{1}{\sigma^{3}}C^{\infty}(S)_{2}[x]$.
Remark 2.
According to [2], it is shown that
$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0.$ (2)
Furthermore, it follows from that
$\mathbb{E}[A(x)]$ $=$ $A^{1}\mathbb{E}[\partial_{1}l(x,\theta)]+A^{2}\mathbb{E}[\partial_{2}l(x,\theta)].$
using (2) we obtain
$\mathbb{E}[A(x)]$ $=$ $0$
[…]
$c$ $=$ $-a(\mu^{2}+\sigma^{2})-b\mu$
By setting $\xi_{2}=\mu^{2}+\sigma^{2}$ et $\xi_{1}=\mu$. $\xi=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$.
So $\xi=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$ is a coordinate system parameterized by $(\mu,\sigma)$. For it suffices show that $\theta$ is a diffeomorphism. Consider the application
$\theta:\mathbb{R}^{2}$ $\rightarrow$ $\mathbb{R}^{2}:(\mu,\sigma)\mapsto(\theta_{1}(\mu,\sigma),\theta_{2}(\mu,\sigma))=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$
by noting that the components $\theta_{1}(\mu,\sigma)$ and $\theta_{2}(\mu,\sigma)$ are differentiable, it is sufficient to show that the Jacobian of $\theta$ is non-zero. Let $J_{M}=\left(\begin{array}[]{cc}\frac{\partial\theta_{1}}{\partial\mu}&\frac{\partial\theta_{1}}{\partial\sigma}\\ \frac{\partial\theta_{2}}{\partial\mu}&\frac{\partial\sigma_{2}}{\partial\sigma}\\ \end{array}\right)$
The Jacobian matrix of $\theta$. It follows that $J_{M}=\left(\begin{array}[]{cc}1&0\\ 2\mu&2\sigma\\ \end{array}\right)$ and its determinant which is the Jacobian is $2\sigma\neq 0$ because $\sigma>0$. It follows that according to the local inversion theorem is locally a diffeomorphism of class $C^{1}$. In coordinates we will need the natural basis and in the system of $\xi$- coordinates we will need the basis associated with the $l$- representation to be determined. To do this we use the Jacobian matrix of the basis passage $\{\mathcal{B}_{\theta}\}$ to the $\xi$-coordinate basis denoted by $\{\mathcal{B}_{\xi}\}=\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}$. Denote by
$(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}\frac{\partial\theta_{1}}{\partial\xi_{1}}&\frac{\partial\theta_{1}}{\partial\xi_{2}}\\ \frac{\partial\theta_{2}}{\partial\xi_{1}}&\frac{\partial\theta_{2}}{\partial\xi_{2}}\\ \end{array}\right)=J_{M}^{-1}$
the inverse of the Jacobian matrix $J_{M}$, and that
$(\partial_{i^{\prime}})=(\overline{\mathcal{B}}^{i}_{\alpha})(\partial_{i})$
. Furthermore, the inverse matrix
$J_{M}^{-1}=\left(\begin{array}[]{cc}1&0\\ \frac{-\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)$
So, the $\xi$- coordinates system is given by $\xi=(\xi_{1},\xi_{2})=(\mu,\mu^{2}+\sigma^{2})$, and we have the $l$- representation in the $\xi$-coordinates system is given by
$\mathcal{B}_{\xi}l(x,\xi)=\{\partial_{\xi_{1}}l(x,\xi),\partial_{\xi_{2}}l(x,\xi)\}$
where $\mathcal{B}_{\xi}=\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}$ is the natural basis of the Amari associated manifold. where $\partial_{\xi_{1}}=\frac{\partial}{\partial\xi_{1}}$ and $\partial_{\xi_{2}}=\frac{\partial}{\partial\xi_{2}}$ thus  $\left\{\frac{\partial}{\partial\xi_{1}},\frac{\partial}{\partial\xi_{2}}\right\}=\mathcal{B}_{\xi}.$
Theorem 1.
Let $(S,g)$ the Riemannian manifold. The natural basis associated to $l$-representation in the coordinated system $(\xi_{1},\xi_{2})$ is given by
$\mathcal{B}(l,\xi)=\left\{\frac{x-\mu}{\sigma^{2}};\frac{(x-\mu)^{2}}{2\sigma^{4}}-\frac{\mu(x-\mu)}{\sigma^{3}}-\frac{1}{2\sigma^{2}}\right\}.$
Proof.
$(\partial_{\xi})=(\bar{\mathcal{B}}^{i}_{\alpha})(\partial_{i})$, $\alpha,i\in\left\{1,2\right\}$, we have
$\left(\begin{array}[]{c}\partial_{\xi_{1}}\\ \partial_{\xi_{2}}\\ \end{array}\right)=\left(\begin{array}[]{cc}1&0\\ -\frac{\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)\left(\begin{array}[]{c}\partial_{1}\\ \partial_{2}\\ \end{array}\right)$
So we have $\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}=\{\partial_{1},\frac{-\mu}{\sigma}\partial_{1}+\frac{1}{2\sigma}\partial_{2}\}$ in $\theta$- coordinates. Its $l$-representation is $\partial_{\xi}l(x,\theta)\}=\left\{\partial_{1}l(x,\theta),-\frac{\mu}{\sigma}\partial_{1}l(x,\theta)+\frac{1}{2\sigma}\partial_{2}l(x,\theta)\right\}$
[…]
Remark 3.
According to [2], the riemannian metric is given by
$g_{ij}(\theta)=<\partial_{i}l(x,\theta),\partial_{j}l(x,\theta)>\textrm{where},g_{ij}(\theta)=\mathbb{E}(\partial_{i}l(x,\theta).\partial_{j}l(x,\theta))=-\mathbb{E}(\partial_{i}\partial_{j}l(x,\theta)).$ (3)
and $g^{ij}$ will be its reverse called stamp information of Fisher. In the same, the metric one associated in the system from the $\xi$- coordinates is defined by
$g_{\alpha\beta}(\xi)=<\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}}>;\;\alpha,\beta\in\{1,2\}$ (4)
2.3 Affine connection.
The collection of all $T_{p}S$ defined a $2n$-dimensional manifold $T(S)$ called Tangent bundle of $S$. More explicitly, $T(S)$ is completely defined by the projection mapping $\pi:TS\longrightarrow S$ such that $\pi^{-1}(\{p\})\cong\mathbb{R}^{n}$ and $\pi$ have the usual local trivialization property. $A:S\longrightarrow TS$ is a vector field on $S$ if $\pi\circ A=1_{S}$. The collection $\mathfrak{X}(S)$ of all vector fields on $S$ is an $C^{\infty}(S)$-module. An affine connection is an $\mathbb{R}$-bilinear mapping
$\nabla:\mathfrak{X}(S)\longrightarrow\mathfrak{X}(S):(A,B)\mapsto\nabla_{A}B$ such that:
$\nabla_{A}(fB)=(Af)(B)+f\nabla_{A}(B),\;\nabla_{fA}(B)=f\nabla_{A}(B).$
In local coordinate $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$ of $S$ , $\nabla_{\partial_{i}}(\partial_{j})=\sum\Gamma_{ij}^{k}\partial_{i}$ and $\Gamma_{ij}^{k}$ is Christoffel coefficient given by: $\Gamma_{ij}^{k}=\frac{1}{2}g^{km}[\partial_{i}.g_{jm}+\partial_{j}.g_{ik}-\partial_{k}.g_{ij}].$
2.4 Torsion tensor and curvature of affine connection.
Let $\nabla$ be an affine-connection on $S$. A torsion tensor of $\nabla$ is an $\mathbb{R}$-bilinear mapping
$T:\mathfrak{X}(S)\times\mathfrak{X}(S)$ $\longrightarrow$ $\mathfrak{X}(S):(A,B)\longmapsto T(A,B)=\nabla_{A}(B)-\nabla_{B}(A)-[A,B]$
[…]
$\mathbb{E}[\partial_{i}l(x,\theta)]$ $=$ $0.$ (5)
[…]
$\Gamma_{ijk}:=\mathbb{E}\left[\partial_{i}\partial_{j}l(x,\theta).\partial_{k}l(x,\theta)\right].$ (6)
The family $\Gamma_{ijk}$ are called $l$-representation of coefficients of the connection $\nabla$ on $S$. With respect to the $l$-representation, $T_{ijk}(\theta)=\langle T(\partial_{\theta_{i}},\partial_{j}),\partial_{k}\rangle,$ it is the coefficient of $T$ with respect to $\partial_{\theta}l(x,\theta)$. Since $T(\partial_{i},\partial_{j})=\nabla_{\partial_{i}}(\partial_{j})-\nabla_{\partial_{j}}(\partial_{i})-(\partial_{i}\partial_{j}-\partial_{j}\partial_{i})$, then $T_{ijk}(\theta)=<\nabla_{\partial_{i}}(\partial_{j})-\nabla_{\partial_{j}}(\partial_{i})-(\partial_{i}\partial_{j}-\partial_{j}\partial_{i}),\partial_{k}>.$ We obtain the following expression: $T_{ijk}(\theta)=<\nabla_{\partial_{i}}(\partial_{j}),\partial_{k}>-<\nabla_{\partial_{j}}(\partial_{i}),\partial_{k}>.$ Therefore,
$T_{ijk}(\theta)=\Gamma_{ijk}(\theta)-\Gamma_{jik}(\theta).$ (7)
The curvature tensor of affine connection $\nabla$ on a $n$-manifold $S$ is a $\mathbb{R}$-trilinear mapping $R:\mathfrak{X}(S)\times\mathfrak{X}(S)\times\mathfrak{X}(S)\longrightarrow\mathfrak{X}(S):(X,Y,Z)\longmapsto R(X,Y)Z$
such that $R(X,Y)Z=[\nabla_{X},\nabla_{Y}]Z-\nabla_{[X,Y]}Z,$ with $\nabla_{X},\nabla_{Y}]Z=\nabla_{X}\nabla_{Y}Z-\nabla_{Y}\nabla_{X}Z.$ In a locale coordinates $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$ of $S$, we have $R(\partial_{i},\partial_{j}).\partial_{k}=\sum_{m=1}^{n}R_{ijk}^{m}\partial_{m}$
[…]
$R_{ijkm}:=\langle R(\partial_{i},\partial_{j},\partial_{k}),\partial_{m}\rangle$ (8)
is the $l$- representation of Riemann-Christoffel curvature tensor of the affine connection $\nabla$ on $S.$ It follow From (8) that $R_{ijkm}$ the following expression:
$R_{ijkm}=\left(\partial_{i}\Gamma^{s}_{jk}-\partial_{j}\Gamma^{s}_{ik}\right)g_{sm}+\left(\Gamma_{irm}\Gamma^{r}_{jk}-\Gamma_{jrm}\Gamma^{r}_{ik}\right).$ (9)
where $g_{sm}$ are coefficients of the Riemannian metric $g$ on $S$. A space with an affine connection is said to be flat when the Riemann-Christoffel curvature and torsion vanishes identically, i.e; $R_{ijkm}(\theta)=0$ holds for any $\theta$. In the coordinate system $\theta=\left(\theta_{1},\dots,\theta_{n}\right)$, the scalar curvature $\mathcal{K}$ of affine connection $\nabla$ is defined by
$\mathcal{K}=\frac{1}{n(n-1)}R_{ijkm}g^{im}g^{jk}.$ (10)
So we have the scalar curve in the frame of reference $\xi$ the noted number $\mathcal{K}$ define by
$\mathcal{K}=\frac{1}{2}R_{\alpha\beta\gamma\eta}g^{\alpha\eta}g^{\beta\gamma}.$ (11)
According to [2], and using 3, we have
$g_{11}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{1}\partial_{1}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[-\frac{1}{\sigma^{2}}\right]$
$=$ $\frac{1}{\sigma^{2}}.$
$g_{22}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{2}\partial_{2}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[\frac{1}{\sigma^{2}}-\frac{3(x-\mu)^{2}}{\sigma^{4}}\right]$
$=$ $-\mathbb{E}\left(\frac{1}{\sigma^{2}}\right)-\mathbb{E}\left[-\frac{3(x-\mu)^{2}}{\sigma^{4}}\right]$
$=$ $-\frac{1}{\sigma^{2}}+3\mathbb{E}\left[\frac{(x-\mu)^{2}}{\sigma^{4}}\right]$
$=$ $-\frac{1}{\sigma^{2}}+3\mathbb{E}\left[\partial_{1}l(x,\theta).\partial_{1}l(x,\theta)\right]$
$=$ $-\frac{1}{\sigma^{2}}-3\mathbb{E}\left[\partial_{1}\partial_{1}l(x,\theta)\right]$
$=$ $-\frac{1}{\sigma^{2}}-3\mathbb{E}\left[-\frac{1}{\sigma^{2}}\right]$
$=$ $-\frac{1}{\sigma^{2}}+\frac{3}{\sigma^{2}}$
$=$ $\frac{2}{\sigma^{2}}$
$g_{12}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{1}\partial_{2}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[\frac{2}{\sigma^{3}}(x-\mu)\right]$
$=$ $-\frac{2}{\sigma^{\sigma}}\mathbb{E}\left[\frac{(x-\mu)}{\sigma^{2}}\right]$
$=$ $-\frac{2}{\sigma}\mathbb{E}\left[\partial_{1}l(x,\theta)\right]$
$=$ $0$
$g_{21}(\theta)$ $=$ $-\mathbb{E}\left[\partial_{2}\partial_{1}l(x,\theta)\right]$
$=$ $-\mathbb{E}\left[-\frac{2}{\sigma^{3}}(x-\mu)\right]$
$=$ $\frac{2}{\sigma^{3}}\mathbb{E}\left[(x-\mu)\right]$
$=$ $\frac{2}{\sigma}\mathbb{E}\left[\frac{(x-\mu)}{\sigma^{2}}\right]$
$=$ $\frac{2}{\sigma}\mathbb{E}\left[\partial_{1}l(x,\theta)\right]$
$=$ $0$
So, the metric one associated for a manifold of gaussian model $S$ is given by
$g_{ij}(\theta)=\left(\begin{array}[]{cc}\frac{1}{\sigma^{2}}&0\\ 0&\frac{2}{\sigma^{2}}\\ \end{array}\right)$ (12)
3 Riemannian metric in $\xi$-coordinates system.
Theorem 2.
Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ be the dual coordinate system associated with the Gaussian statistical manifold. Let $\left(g_{ij}\right)_{1\leq i,j\leq 2}$ denote the Riemannian metric expressed in the natural coordinate system $\theta=(\mu,\sigma)$, as defined by equation (12). Let $(\overline{\mathcal{B}}^{i}_{\alpha})$ denote the inverse of the Jacobian matrix of the coordinate transformation from $\theta$ to $\xi$, given explicitly by: $(\overline{\mathcal{B}}^{i}_{\alpha})=\left(\begin{array}[]{cc}1&0\\ \frac{-\mu}{\sigma}&\frac{1}{2\sigma}\\ \end{array}\right)$ Then, the components of the Riemannian metric $g_{\alpha\beta}(\xi)$ in the $\xi$-coordinate system are given by:
$g_{\alpha\beta}(\xi)=\frac{1}{\sigma^{2}}\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}+\frac{2}{\sigma^{2}}\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}$
$\alpha,\;\beta\in\left\{1,2\right\}$.
Proof.
Let $\theta=(\mu,\sigma)$ denote the natural coordinate system and $\xi=(\mu,\mu^{2}+\sigma^{2})$ the dual coordinate system on the Gaussian manifold. From the coordinate transformation, the Jacobian matrix of the map $\theta\mapsto\xi$ is given by:
$J=\begin{pmatrix}\frac{\partial\xi_{1}}{\partial\mu}&\frac{\partial\xi_{1}}{\partial\sigma}\\ \frac{\partial\xi_{2}}{\partial\mu}&\frac{\partial\xi_{2}}{\partial\sigma}\end{pmatrix}=\begin{pmatrix}1&0\\ 2\mu&2\sigma\end{pmatrix}$,
whose inverse is:
$J^{-1}=(\overline{\mathcal{B}}^{i}_{\alpha})=\begin{pmatrix}1&0\\ -\frac{\mu}{\sigma}&\frac{1}{2\sigma}\end{pmatrix}$.
Let $\{\partial_{1},\partial_{2}\}=\left\{\frac{\partial}{\partial\mu},\frac{\partial}{\partial\sigma}\right\}$ be the basis associated with the $\theta$-coordinates, and $\{\partial_{\xi_{1}},\partial_{\xi_{2}}\}$ the basis in the $\xi$-coordinate system. Using the chain rule and the inverse Jacobian, the basis transformation is
$\partial_{\xi_{\alpha}}=\sum\limits_{i=1}^{2}\bar{B_{\alpha}}^{i}\partial_{i}=\bar{B_{\alpha}}^{1}\partial_{1}+\bar{B_{\alpha}}^{2}\partial_{2},\;and\;\partial_{\xi_{\beta}}=\sum\limits_{j=1}^{2}\bar{B_{\beta}}^{j}\partial_{j}=\bar{B_{\beta}}^{1}\partial_{1}+\bar{B_{\beta}}^{2}\partial_{2}.$
Then, the components of the Riemannian metric in the $\xi$-coordinates are given by the inner product
$g_{\alpha\beta}(\xi)=<\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}}>,\;\alpha,\;\beta\in\left\{1,2\right\}.$
or $\bar{B_{\alpha}}^{i}=\left(\frac{\partial\theta_{i}}{\partial\xi_{\alpha}}\right)$ We have
$g_{\alpha\beta}(\xi)$ $=$ $<\bar{B_{\alpha}}^{1}\partial_{1}+\bar{B_{\alpha}}^{2}\partial_{2},\bar{B_{\beta}}^{1}\partial_{1}+\bar{B_{\beta}}^{2}\partial_{2}>$
$=$ $<\bar{B_{\alpha}}^{1}\partial_{1},\bar{B_{\beta}}^{1}\partial_{1}>+<\bar{B_{\alpha}}^{1}\partial_{1},\bar{B_{\beta}}^{2}\partial_{2}>+<\bar{B_{\alpha}}^{2}\partial_{2},\bar{B_{\beta}}^{1}\partial_{1}>+<\bar{B_{\alpha}}^{2}\partial_{2},\bar{B_{\beta}}^{2}\partial_{2}>$
$=$ $\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}g_{11}(\theta)+\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}g_{22}(\theta)$
$\forall\alpha,\;\beta\in\left\{1,2\right\}$. So,
$g_{\alpha\beta}(\xi)=\frac{1}{\sigma^{2}}\bar{B_{\alpha}}^{1}\bar{B_{\beta}}^{1}+\frac{2}{\sigma^{2}}\bar{B_{\alpha}}^{2}\bar{B_{\beta}}^{2}$
While being (12) and (2) there is the proof of following theorem
Theorem 3.
Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ be the dual coordinate system associated with the natural coordinates $\theta=\left(\mu,\sigma\right)$ on the Gaussian statistical manifold. Then, the Riemannian metric $G_{d}$ in the $\xi$-coordinate system is given by:
$G_{d}=\frac{1}{\sigma^{4}}\left(\begin{array}[]{cc}\sigma^{2}+2\mu^{2}&-\mu\\ -\mu&\frac{1}{2}\end{array}\right),$
The determinant of this metric tensor is:
$DetG_{d}=\frac{1}{2\sigma^{6}}$
The inverse of the Fisher information matrix in the $\xi$-coordinate system, denoted $G_{d}^{-1}$, is given by:
$G_{d}^{-1}=\left(\begin{array}[]{cc}\sigma^{2}&2\mu\sigma^{2}\\ 2\mu\sigma^{2}&2\sigma^{4}+4\mu^{2}\sigma^{2}\end{array}\right)$
Proof.
Using theorem 2, we have
$g_{11}(\xi)=\bar{B_{1}}^{1}\bar{B_{1}}^{1}g_{11}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{2}g_{22}(\theta)$
$g_{11}(\xi)=\left(\frac{\partial\theta_{1}}{\partial\xi_{1}}\right)\left(\frac{1}{\sigma^{2}}\right)+\left(\frac{\partial\theta_{2}}{\partial\xi_{1}}\right)\left(\frac{2}{\sigma^{2}}\right)$
because $\theta_{1}=\mu,\;\xi_{1}=\mu$
$g_{11}(\xi)$ $=$ $\frac{1}{\sigma^{2}}+\frac{2\mu^{2}}{\sigma^{4}}$
$=$ $\frac{1}{\sigma^{2}}\left(\frac{\sigma^{2}+2\mu^{2}}{\sigma^{2}}\right)$
$=$ $\frac{1}{\sigma^{4}}\left(\sigma^{2}+2\mu^{2}\right)$
then we have
$g_{22}(\xi)$ $=$ $\bar{B_{2}}^{1}\bar{B_{2}}^{1}g_{11}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}g_{22}(\theta)$
$g_{22}(\xi)$ $=$ $\left(\bar{B_{2}}^{1}\right)^{2}g_{11}(\theta)+\left(\bar{B_{2}}^{2}\right)^{2}g_{22}(\theta)$
$g_{22}(\xi)$ $=$ $\left(\bar{B_{2}}^{1}\right)^{2}\left(\frac{1}{\sigma^{2}}\right)+\left(\bar{B_{2}}^{2}\right)^{2}\left(\frac{2}{\sigma^{2}}\right)$
$g_{22}(\xi)$ $=$ $\left(\frac{1}{2\sigma}\right)^{2}\left(\frac{2}{\sigma^{2}}\right)$
$=$ $\frac{2}{\sigma^{4}}$
$i.e.\;g_{22}(\xi)$ $=\frac{1}{2\sigma^{4}}$
$g_{12}(\xi)$ $=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{1}g_{11}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}g_{22}(\theta)$
$=$ $\frac{-\mu}{\sigma^{4}}$
$g_{21}(\xi)$ $=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}g_{11}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{2}g_{12}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{1}g_{21}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}g_{22}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}g_{11}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}g_{22}(\theta)$
$=$ $\frac{-\mu}{\sigma^{4}}$
we have the following metric
$G_{d}=\frac{1}{\sigma^{4}}\left(\begin{array}[]{cc}\sigma^{2}+2\mu^{2}&-\mu\\ -\mu&\frac{1}{2}\end{array}\right),$
and
$DetG_{d}=\frac{1}{2\sigma^{6}}$
So, we have the result ∎
4 Torsion and curvature tensor in $\xi$-coordinates system.
Theorem 4.
Let $\nabla^{(1)}$ be the $\alpha$-connection of Amari with $\alpha=1$. Then, the torsion tensor $T^{\gamma}_{\alpha\beta}(\xi)$ in the dual coordinate system $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ is non-zero. This is in contrast to the natural coordinate system $\theta$, where the torsion vanishes for the Levi-Civita connection. The presence of torsion in $\xi$-coordinates reflects the fact that duality does not preserve symmetry of the connection.
Proof.
According to Amari’s [2], we have Christoffel’s coefficient in $\theta$-coordinate system on gaussian statistical manifold
$\Gamma^{(1)}_{111}(\theta)=\Gamma^{(1)}_{112}(\theta)=\Gamma^{(1)}_{221}(\theta)=\Gamma^{(1)}_{122}(\theta)=\Gamma^{(1)}_{212}(\theta)=0$
$\Gamma^{(1)}_{121}(\theta)=\Gamma^{(1)}_{211}(\theta)=-\frac{2}{\sigma^{3}},\;\Gamma^{(1)}_{222}(\theta)=-\frac{6}{\sigma^{3}}.$
Calculation of the coefficients of christoffel’s in the system of coordinates $\xi$.
$\Gamma^{1}_{111}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{1}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$+\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $-\frac{\mu}{\sigma}.(\frac{-2}{\sigma^{3}})-\frac{\mu}{\sigma}.(\frac{-2}{\sigma^{3}})+\frac{-\mu^{3}}{\sigma^{3}}.(\frac{-6}{\sigma^{3}})$
$=$ $\frac{4\mu\sigma^{2}+6\mu^{3}}{\sigma^{6}}$
$\Gamma_{111}(\xi)$ $=$ $\frac{4\mu\sigma^{2}+6\mu^{3}}{\sigma^{6}}$
By analogy with calculation preceding one a:
$\Gamma_{112}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{1}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{1}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\frac{\mu^{2}}{\sigma^{2}}.(\frac{1}{2\sigma})(\frac{-6}{\sigma^{3}})$
$\Gamma_{112}(\xi)$ $=$ $-\frac{3\mu^{2}}{\sigma^{6}}$
In the same way,
$\Gamma_{221}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{2}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{221}(\xi)$ $=$ $\frac{3\mu}{2\sigma^{6}}$
$\Gamma_{212}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{1}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{212}(\xi)$ $=$ $\frac{3\mu}{2\sigma^{6}}$
In the same way in a similar way,
$\Gamma_{121}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{2}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$\ \Gamma_{121}(\xi)$ $=$ $-\frac{1}{\sigma^{4}}-\frac{3\mu^{2}}{\sigma^{6}}$
In the same way one obtains
$\Gamma_{211}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{1}}^{j}\bar{B_{1}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{1}}^{1}\bar{B_{1}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{1}}^{2}\bar{B_{1}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{211}(\xi)$ $=$ $-\frac{1}{\sigma^{4}}-\frac{3\mu^{2}}{\sigma^{6}}$
$\Gamma_{122}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{1}}^{i}\bar{B_{2}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{1}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{1}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{122}(\xi)$ $=$ $\frac{3\mu}{2\sigma^{6}}$
$\Gamma_{222}(\xi)$ $=$ $\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{2}\sum\limits_{k=1}^{2}\bar{B_{2}}^{i}\bar{B_{2}}^{j}\bar{B_{2}}^{k}\Gamma_{ijk}(\theta)$
$=$ $\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{111}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{112}(\theta)+$
$\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{121}(\theta)+\bar{B_{2}}^{1}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{122}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{1}\Gamma_{211}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+$
$\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{1}\Gamma_{221}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{1}\bar{B_{2}}^{2}\Gamma_{212}(\theta)+\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$=$ $\bar{B_{2}}^{2}\bar{B_{2}}^{2}\bar{B_{2}}^{2}\Gamma_{222}(\theta)$
$\Gamma_{222}(\xi)$ $=$ $-\frac{3}{4\sigma^{6}}$
Using 7, we have the result. ∎
Theorem 5.
Let $\xi=\left(\mu,\mu^{2}+\sigma^{2}\right)$ be the dual coordinate system associated with the Gaussian statistical manifold. Then, the components of the Riemann curvature tensor of the affine connection $\nabla^{(1)}$ in these coordinates include:
$R_{1111}(\xi)$ $=$ $R_{2222}(\xi)=R_{1122}(\xi)=R_{1112}(\xi)=R_{2121}(\xi)=R_{2211}(\xi)=0$
$R_{2221}(\xi)=R_{2212}(\xi)=R_{2122}(\xi)=R_{1121}(\xi)=R_{1211}(\xi)=0$
$R_{1221}(\xi)$ $=$ $\frac{\mu}{\sigma^{6}},\;R_{2112}(\xi)=\frac{6\sigma^{8}+3\mu\sigma^{6}+6\mu^{2}\sigma^{6}+6\mu^{2}}{4\sigma^{14}},\;R_{1212}(\xi)=\frac{1}{2\sigma^{6}}$
$R_{2111}(\xi)$ $=$ $\frac{-2\sigma^{8}+2\mu^{2}\sigma^{6}+46\mu^{2}\sigma^{8}+24\mu^{4}\sigma^{6}+12\mu^{2}\sigma^{2}-4\mu\sigma^{2}+6\sigma^{10}+12\mu^{4}-3\mu^{2}+9\mu\sigma^{8}+18\mu^{3}\sigma^{6}}{2\sigma^{14}}$
and the scalar curve is given by
$\mathcal{K}$ $=$ $\frac{16\mu^{2}\sigma^{8}+12\mu^{4}+6\sigma^{10}-\mu\sigma^{8}-2\mu^{2}\sigma^{2}+10\mu^{3}\sigma^{6}+92\mu^{3}\sigma^{8}+48\mu^{5}\sigma^{6}}{4\sigma^{10}}$
$+\frac{12\mu\sigma^{10}+24\mu^{5}-6\mu^{3}}{4\sigma^{10}}$
Proof.
According to Amari [2], we have
$\Gamma_{\alpha\beta}^{\gamma}(\xi)=\frac{1}{2}g^{\gamma\lambda}(\xi)\left(\frac{\partial g_{\beta\lambda}(\xi)}{\partial\xi_{\alpha}}+\frac{\partial g_{\gamma\alpha}(\xi)}{\partial\xi_{\beta}}-\frac{\partial g_{\alpha\beta}(\xi)}{\partial\xi_{\gamma}}\right)$
with $\alpha,\;\beta,\;\gamma,\;\lambda\in\left\{1,2\right\}$. So, we have the coefficients following
$\Gamma_{11}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}$
$\Gamma_{11}^{1}(\xi)$ $=$ $\frac{\mu}{\sigma^{2}}.$
$\Gamma_{22}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}+\frac{1}{2}g^{22}(\xi)\frac{\partial g_{22}(\xi)}{\partial\xi_{2}}$
$\Gamma_{22}^{2}(\xi)$ $=$ $0.$
$\Gamma_{12}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}\right)+\frac{1}{2}g^{22}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}\right)$
$\Gamma_{12}^{2}(\xi)$ $=$ $-\frac{\mu}{\sigma^{2}}.$
$\Gamma_{21}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}\right)+\frac{1}{2}g^{22}(\xi)\left(\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}\right)$
$\Gamma_{21}^{2}(\xi)$ $=$ $\frac{\mu}{\sigma^{2}}.$
$\Gamma_{11}^{2}(\xi)$ $=$ $\frac{1}{2}g^{21}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}\right)+\frac{1}{2}g^{22}(\xi)\left(\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}\right)$
$\Gamma_{11}^{2}(\xi)$ $=$ $\frac{-2\mu^{2}+\mu-3\sigma^{8}-6\mu^{2}\sigma^{6}}{\sigma^{8}}.$
$\Gamma_{21}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}}{\partial\xi_{1}}-\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}\right)$
$\Gamma_{21}^{1}(\xi)$ $=$ $\frac{1}{2\sigma^{2}}.$
$\Gamma_{22}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{2}}+\frac{\partial g_{12}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}\right)$
$\Gamma_{22}^{1}(\xi)$ $=$ $0.$
$\Gamma_{12}^{1}(\xi)$ $=$ $\frac{1}{2}g^{11}(\xi)\left(\frac{\partial g_{21}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{12}(\xi)}{\partial\xi_{1}}\right)+\frac{1}{2}g^{12}(\xi)\left(\frac{\partial g_{22}(\xi)}{\partial\xi_{1}}+\frac{\partial g_{11}(\xi)}{\partial\xi_{2}}-\frac{\partial g_{1^{\prime}2^{\prime}}}{\partial\xi^{1}}\right)$
$\Gamma_{12}^{1}(\xi)$ $=$ $8\mu^{2}\sigma^{4}+\frac{4\mu+1}{2\sigma^{2}}.$
According to Amari [2], the tensor of Riemanienne curvature of Christoffel is given by
$R_{\alpha\beta\gamma\eta}$ $=$ $<R(\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}},\partial_{\xi_{\gamma}}),\partial_{\xi_{\eta}}>$ (13)
we have
$R(\partial_{\xi_{\alpha}},\partial_{\xi_{\beta}},\partial_{\xi_{\gamma}})$ $=$ $\left(\nabla_{\partial_{\xi_{\alpha}}}\nabla_{\partial_{\xi_{\beta}}}-\nabla_{\partial_{\xi_{\beta}}}\nabla_{\partial_{\xi_{\alpha}}}\right)\partial_{\xi_{\alpha}}-\nabla_{(\partial_{\xi_{\alpha}}\partial_{\xi_{\beta}}-\partial_{\xi_{\beta}}\partial_{\xi_{\alpha}})}\partial_{\xi_{\gamma}}$
$=$ $\left(\nabla_{\partial_{\xi_{\alpha}}}\nabla_{\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})-\nabla_{\partial_{\xi_{\beta}}}\nabla_{\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})\right)-\nabla_{\partial_{\xi_{\alpha}}\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})+\nabla_{\partial_{\xi_{\beta}}\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})$
$=$ $\left(\nabla_{\partial_{\xi_{\alpha}}}\Gamma_{\beta\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}-\nabla_{\partial_{\xi_{\beta}}}\Gamma_{\alpha\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}\right)-\partial_{\alpha}\nabla_{\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})+\partial_{\xi_{\beta}}\nabla_{\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})$
$=$ $\left(\partial_{\xi_{\alpha}}\Gamma_{\beta\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\beta\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\alpha}}}\partial_{\xi_{s^{\prime}}}\right)-\left(\partial_{\xi_{\beta}}\Gamma_{\alpha\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\alpha\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\beta}}}\partial_{\xi_{s^{\prime}}}\right)$
$-\partial_{\xi_{\alpha}}\nabla_{\partial_{\xi_{\beta}}}(\partial_{\xi_{\gamma}})+\partial_{\xi_{\beta}}\nabla_{\partial_{\xi_{\alpha}}}(\partial_{\xi_{\gamma}})$
$=$ $\left(\partial_{\xi_{\alpha}}\Gamma_{\beta\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\beta\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\alpha}}}\partial_{\xi_{s^{\prime}}}\right)-\left(\partial_{\xi_{\beta}}\Gamma_{\alpha\gamma}^{s^{\prime}}\partial_{\xi_{s^{\prime}}}+\Gamma_{\alpha\gamma}^{s^{\prime}}\nabla_{\partial_{\xi_{\beta}}}\partial_{\xi_{s^{\prime}}}\right)$
$-\partial_{\xi_{\alpha}}\Gamma_{\beta\gamma}^{r^{\prime}}\partial_{\xi_{r^{\prime}}}+\partial_{\xi_{\beta}}\Gamma_{\alpha\gamma}^{r^{\prime}}\partial_{\xi_{r^{\prime}}}$
While replacing in (13) we obtain
$R_{\alpha\beta\gamma\eta}$ $=$ $\left(\partial_{\xi_{\alpha}}\Gamma^{s^{\prime}}_{\beta\gamma}-\partial_{\xi_{\beta}}\Gamma^{s^{\prime}}_{\alpha\gamma}\right)g_{s^{\prime}\eta}-\left(\Gamma^{r^{\prime}}_{\beta\gamma}\Gamma_{\alpha r^{\prime}\eta}-\Gamma^{r^{\prime}}_{\alpha\gamma}\Gamma_{\beta r^{\prime}\eta}\right)$
$+<\Gamma^{s^{\prime}}_{\beta\gamma}\nabla_{\partial_{\xi_{\alpha}}}\partial_{\xi_{s^{\prime}}},\partial_{\xi_{\eta}}>-<\Gamma^{s^{\prime}}_{\alpha\gamma}\nabla_{\partial_{\xi_{\beta}}}\partial_{\xi_{s^{\prime}}},\partial_{\xi_{\eta}}>$
with $\alpha,\beta,\gamma,\eta,s^{\prime}\in\left\{1,2\right\}$ so,
$R_{1111}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{111}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{121}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{121}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>+<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>-<\Gamma_{11}^{2}\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>$
$=$ $0$
$R_{2222}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{222}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{222}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{222}(\xi)\Gamma_{22}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{22}^{2}(\xi)\right)$
$+<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ $0$
$R_{1122}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{1^{\prime}}\Gamma_{12}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{112}(\theta)\Gamma_{12}^{1}(\xi)\right)-\left(\Gamma_{122}(\xi)\Gamma_{12}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{12}^{2}(\xi)\right)$
$+<\Gamma_{12}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{12}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{12}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{12}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ $0.$
$R_{1112}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)-\partial_{1^{\prime}}\Gamma_{11}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{122}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ $0$
$R_{2121}(\xi)$ $=$ $\left(\partial_{2^{\prime}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{111}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{12}^{2}(\xi)-\Gamma_{121}(\xi)\Gamma_{22}^{2}(\xi)\right)$
$+<\Gamma_{12}^{1}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>+<\Gamma_{12}^{2}(\xi)\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$-<\Gamma_{22}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>-<\Gamma_{22}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$=$ $0$
$R_{2211}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{211}\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $0.$
$R_{2221}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{22}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)\right)$
$+\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)+\Gamma_{221}\Gamma_{22}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{22}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{22}^{2}(\xi)$
$=$ $0.$
$R_{2212}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{212}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{212}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{222}\Gamma_{21}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{211}\Gamma_{21}^{1}(\xi)+\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}\Gamma_{21}^{1}(\xi)-\Gamma_{221}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $0.$
$R_{2122}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{212}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{22}^{1}(\xi)\right)-\left(\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{22}^{2}(\xi)\right)$
$+\Gamma_{212}(\xi)\Gamma_{12}^{1}(\xi)+\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}\Gamma_{22}^{1}(\xi)-\Gamma_{122}(\xi)\Gamma_{22}^{2}(\xi)$
$=$ $0.$
$R_{1221}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{22}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{12}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{22}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{12}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}\Gamma_{22}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{121}(\xi)\Gamma_{22}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{12}^{2}(\xi)\right)$
$+\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{221}\Gamma_{12}^{2}(\xi)$
$=$ $\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $\frac{\mu}{\sigma^{6}}.$
$R_{2112}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{21}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{212}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{212}(\xi)\Gamma_{11}^{1}(\xi)+\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)$
$=$ $\frac{6\sigma^{8}+3\mu\sigma^{6}+6\mu^{2}\sigma^{6}+6\mu^{2}}{4\sigma^{14}}.$
$R_{1212}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{212}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+<\Gamma_{21}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>+<\Gamma_{21}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$-<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{11}^{2}\nabla_{\partial_{\xi_{2}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=\left(\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{1}\right)g_{12}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)\right)g_{22}(\xi)$
$-\left(\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{212}\Gamma_{11}^{1}\right)-\left(\Gamma_{122}\Gamma_{21}^{2}-\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{212}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{222}(\xi)\Gamma_{11}^{2}(\xi)$
$=$ $\frac{1}{2\sigma^{6}}.$
$R_{1211}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{21}^{1}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{21}^{2}(\xi)-\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)-\Gamma_{211}(\xi)\Gamma_{11}^{1}(\xi)\right)-\left(\Gamma_{121}\Gamma_{21}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)\right)$
$+\Gamma_{111}(\xi)\Gamma_{21}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{21}^{2}(\xi)-\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{211}(\xi)\Gamma_{11}^{1}(\xi)$
$=$ $0.$
$R_{1121}(\xi)$ $=$ $\left(\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{12}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)\right)-\left(\Gamma_{121}\Gamma_{12}^{2}(\xi)-\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)\right)$
$+\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)+\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)-\Gamma_{111}(\xi)\Gamma_{12}^{1}(\xi)-\Gamma_{121}(\xi)\Gamma_{12}^{2}(\xi)$
$=$ $0.$
$R_{2111}(\xi)$ $=$ $\left(\partial_{\xi_{2}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}(\xi)\Gamma_{11}^{1}(\xi)-\Gamma_{112}(\xi)\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+<\Gamma_{11}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{1}}>+<\Gamma_{11}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{1}}>$
$-<\Gamma_{21}^{1}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{1}}),\partial_{\xi_{2}}>-<\Gamma_{21}^{2}(\xi)\nabla_{\partial_{\xi_{1}}}(\partial_{\xi_{2}}),\partial_{\xi_{2}}>$
$=$ $\left(\partial_{2^{\prime}}\Gamma_{11}^{1}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{1}(\xi)\right)g_{11}(\xi)+\left(\partial_{\xi_{2}}\Gamma_{11}^{2}(\xi)-\partial_{\xi_{1}}\Gamma_{11}^{2}(\xi)\right)g_{21}(\xi)$
$-\left(\Gamma_{211}\Gamma_{11}^{1}(\xi)-\Gamma_{112}\Gamma_{21}^{1}(\xi)\right)-\left(\Gamma_{221}(\xi)\Gamma_{11}^{2}(\xi)-\Gamma_{122}(\xi)\Gamma_{21}^{2}(\xi)\right)$
$+\Gamma_{11}^{1}(\xi)\Gamma_{111}(\xi)+\Gamma_{11}^{1}(\xi)\Gamma_{121}(\xi)-\Gamma_{21}^{1}(\xi)-\Gamma_{21}^{2}(\xi)\Gamma_{122}(\xi)$
$=$ $\frac{-2\sigma^{8}+2\mu^{2}\sigma^{6}+46\mu^{2}\sigma^{8}+24\mu^{4}\sigma^{6}+12\mu^{2}\sigma^{2}-4\mu\sigma^{2}+6\sigma^{10}+12\mu^{4}-3\mu^{2}+9\mu\sigma^{8}+18\mu^{3}\sigma^{6}}{2\sigma^{14}}$
Using the definition 11, we have
$\mathcal{K}$ $=$ $\frac{1}{2}R_{1111}(\xi)g^{11}(\xi)g^{11}(\xi)+\frac{1}{2}R_{1112}(\xi)g^{12}(\xi)g^{11}(\xi)+\frac{1}{2}R_{1121}(\xi)g^{11}(\xi)g^{12}(\xi)$
$+\frac{1}{2}R_{1122}(\xi)g^{12}(\xi)g^{12}(\xi)+\frac{1}{2}R_{1211}(\xi)g^{11}(\xi)g^{21}(\xi)+\frac{1}{2}R_{1212}(\xi)g^{12}(\xi)g^{21}(\xi)$
$+\frac{1}{2}R_{1221}(\xi)g^{11}(\xi)g^{22}(\xi)+\frac{1}{2}R_{1222}(\xi)g^{12}(\xi)g^{22}(\xi)+\frac{1}{2}R_{2111}(\xi)g^{21}(\xi)g^{11}(\xi)$
$+\frac{1}{2}R_{2112}(\xi)g^{22}(\xi)g^{11}(\xi)+\frac{1}{2}R_{2121}(\xi)g^{21}(\xi)g^{12}(\xi)+\frac{1}{2}R_{2122}(\xi)g^{22}(\xi)g^{12}(\xi)$
$+\frac{1}{2}R_{2211}(\xi)g^{21}(\xi)g^{21}(\xi)+\frac{1}{2}R_{2212}(\xi)g^{22}g^{21}(\xi)+\frac{1}{2}R_{2221}(\xi)g^{21}(\xi)g^{22}(\xi)$
$+\frac{1}{2}R_{2222}(\xi)g^{22}(\xi)g^{22}(\xi)$
$=$ $\frac{1}{2}R_{1212}(\xi)g^{12}(\xi)g^{21}(\xi)+\frac{1}{2}R_{2111}(\xi)g^{21}(\xi)g^{11}(\xi)+\frac{1}{2}R_{2112}(\xi)g^{22}(\xi)g^{11}(\xi)$
$=$ $\frac{16\mu^{2}\sigma^{12}+12\mu^{4}\sigma^{4}+6\sigma^{14}-\mu\sigma^{12}-2\mu^{2}\sigma^{6}+10\mu^{3}\sigma^{10}+92\mu^{3}\sigma^{12}+48\mu^{5}\sigma^{10}}{4\sigma^{14}}$
$+\frac{12\mu\sigma^{14}+24\mu^{5}\sigma^{4}-6\mu^{3}\sigma^{4}}{4\sigma^{14}}$
Thus the curve will be given by
$\mathcal{K}$ $=$ $\frac{16\mu^{2}\sigma^{8}+12\mu^{4}+6\sigma^{10}-\mu\sigma^{8}-2\mu^{2}\sigma^{2}+10\mu^{3}\sigma^{6}+92\mu^{3}\sigma^{8}+48\mu^{5}\sigma^{6}}{4\sigma^{10}}$
$+\frac{12\mu\sigma^{10}+24\mu^{5}-6\mu^{3}}{4\sigma^{10}}$
5 General conclusion
This study has examined the differential geometric structure of the Gaussian statistical model when expressed in a dual coordinate system $\xi=(\mu,\mu^{2}+\sigma^{2})$. Through explicit computations of the Fisher information metric, the affine connection, and the curvature tensors, we have shown that the dual parametrization introduces non-trivial geometric effects. In contrast to the classical coordinate system $\theta=(\mu,\sigma)$, where the manifold is known to be both flat and torsion-free, the $\xi$-system exhibits non zero torsion and scalar curvature. These findings highlight an important phenomenon in information geometry: duality does not guarantee invariance of geometric quantities such as curvature and torsion. In particular, our results show that the choice of coordinate system can deeply influence the global geometric interpretation of a statistical model, even for well understood cases like the normal distribution. From a broader perspective, this work invites further exploration into the stability and variability of geometric invariants under coordinate transformations. Understanding which geometric properties are intrinsic to the model and which depend on parametrization remains a central question. Future research could aim to classify transformations that preserve curvature or to extend the analysis to other statistical families, such as the beta, Weibull, or log-normal distributions. Ultimately, this study reinforces the idea that the geometry of statistical models is not only local and computational it is also deeply sensitive to the way we represent information.
[…]
Supplementary information This manuscript has no additional data.
[…]
Acknowledgments We would like to thank all the active members of the algebra and geometry research group at the University of Maroua in Cameroon.
Declarations
This article has no conflict of interest to the journal. No financing with a third party.
  • No Funding
  • No Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use)
  • Ethics approval
  • Consent to participate
  • Consent for publication
  • Availability of data and materials
  • Code availability
  • Authors’ contributions